OpenStudy (anonymous):

Please help me, I've been having trouble with this question. Find an equation of the line that passes through the point (-2, 2) and is parallel to the line 2x - 4y - 8 = 0 Thank you.

OpenStudy (campbell_st):

rewrite the parallel line in slope intercept form y = 1/2 x - 2 whats the slope of the line...?

OpenStudy (anonymous):

I did that but I don't know what next to do. I plotted the (-2, 2) and I simplified 2x -4y - 8 = 0 to the slope intercept eqation. I don't know what's missing. @campbell_st

OpenStudy (anonymous):

Wouldn't the slope be 1/2?

OpenStudy (campbell_st):

ok so I've written the parallel line equation in slope intercept for above. Can you tell me the slope of the line... you don't need to graph it...

OpenStudy (anonymous):

The slope is 1/2, I hope i'm correct. Otherwise, I do not know.

OpenStudy (campbell_st):

yep.... thats correct... the the slope intercept for of an equation is y = mx + b you know m is 1/2 so its y = 1/2 x + b to find b use the point. Substitute x = 2 and y = 2 into the equation to find b

OpenStudy (anonymous):

That makes sense! Thanks a lot! Really appreciate it :)